OpenStudy (anonymous):

an object is thrown downward with an initial (t=0) speed of 10 m/s from a height 60 m above the ground , at the same instant ( t=0) , a second object is propelled vertically upward from ground level with a speed of 40 m/s . at what height above the ground will the two objects pass each other?

OpenStudy (anonymous):

do u hav answer in back

OpenStudy (anonymous):

10+ 40 = 50 m/s 60/50 =1.2 sec

OpenStudy (anonymous):



OpenStudy (anonymous):

see the rate at which the object which is thrown up loses it speed is same to the rate at which the object thrown downward gains the speed......... the rate of change of speed = g =9.8m/s^2 so relative vel. of one of object with respect to other = v1 + v2 =10+40 = 50 m/s {because the sum of speeds always remains constt.)

OpenStudy (anonymous):

use 2nd eqn of motion to find the distance of meeting in any case either upward or downward taking g=+ve for downward motion and g=-ve for upward motion

OpenStudy (anonymous):

Ok so first I mapped out the known factors in the equation for each motion of interest For motion A we have: X0 = 60, VX0 = 10, AX0 = 9.8, missing Δt and Xf For motion B we have: X0 = 0, VX0 = 40, AX0 = 9.8, again missing Δt and Xf I figure that we are looking for the X when both x's are zero so I set them to equal each other in the same form as the equation AX0 + AVX0(Δt) + (1/2)(AX0)(Δt)^2 = BX0 + BVX0(Δt) + (1/2)(AX0)(Δt)^2 60 + 10t + 4.9t^2 = 40t - 4.9t^2 From this I get everything to one side so the right side is 0 and I can use the quadratic formula to solve for Δt Equation looks like 9.8t^2 - 30t + 60 = 0 So i do the quadratic equation 30 +/- sqrt(30^2 - 4(9.8)(60)) all over 2(9.8) I end up with 3.47475 OR -.413526 as Δt. I plug either one of those numbers into either one of the original equations to solve for Xf and I get numbers that are no where near correct. He has listed the answer as 41 m. Reference https://www.physicsforums.com/threads/where-when-physics-problem-with-1-obj-going-up-and-the-other-going-down.572573/

OpenStudy (anonymous):

the answer says it's 37 but i am not sure about it

OpenStudy (anonymous):

and when i get t how am i going to substitute it in one of the original equations , when i do it i don't get 41 or 37!